OpenStudy (anonymous):

Find a cubic function with the given zeros. -6, 7, -4 @jim_thompson5910 @dan815 @amistre64 @e.mccormick @kainui @sleepyjess @robtobey @eclipsedstar @precal

OpenStudy (precal):

p(x)=(x+6)(x-7)(x+4)

OpenStudy (precal):

recall (x-r) is your factor if you know that x=-3 then your factor form is (x+3)

OpenStudy (anonymous):

okay so what do i do next @precal

OpenStudy (precal):

well that is your answer unless they want you to expand it

OpenStudy (anonymous):

f(x) = x3 + 3x2 + 46x - 168 f(x) = x3 + 3x2 - 46x + 168 f(x) = x3 - 3x2 - 46x - 168 f(x) = x3 + 3x2 - 46x - 168 these are my options @precal

OpenStudy (precal):

expand it out

OpenStudy (precal):

you can do it

OpenStudy (anonymous):

i honestly cant figure it out @precal

OpenStudy (precal):

did you not learn to expand binomials?

OpenStudy (anonymous):

no @precal

OpenStudy (precal):

(x+1)(x+4)

OpenStudy (e.mccormick):

It is multiplication. That is it. Know how to FOIL a pair of binomials?

OpenStudy (precal):

FOIL First Outer Inner Last

OpenStudy (precal):

all you do is distribute x to (X+4) then distribute 1 to (x+4) then combine like terms

OpenStudy (anonymous):

so its x^2+5x+4 right? @precal

OpenStudy (precal):

yes now pick two of the three I gave you before and do the same, I will check your solution

OpenStudy (anonymous):

like solve (x+6)(x-7)? @precal

OpenStudy (precal):

yes you can do it, I will check your solution

OpenStudy (anonymous):

is this the answer? f(x) = x3 + 3x2 - 46x + 168 @precal

OpenStudy (precal):

Let's do it one step at a time what did you get for (x+6)(x-7)

OpenStudy (anonymous):

x^2-x-42 @precal

OpenStudy (precal):

yes now expand (x+4)(x^2-x-42) remember distribute x to (x^2-x-42) and then distribute 4 to (x^2-x-42) collect like terms

OpenStudy (anonymous):

x^3+3x^2-46x-168 @precal

OpenStudy (precal):

you got it !!!! GOOD JOB

OpenStudy (precal):

did you understand the process?

OpenStudy (anonymous):

can you help me with another one? @precal

OpenStudy (precal):

ok what do you have?

OpenStudy (anonymous):

State the domain of the rational function. f(x) = 12/13-x @precal

OpenStudy (precal):

Easy the domain is all real numbers except 13 because 13 creates a zero in the denominator. Recall N/0 for no zero on the bottom Recall O/K Ok to have a zero on the top

OpenStudy (anonymous):

one more? @precal

OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 7, -11, and 2 + 8i f(x) = x4 - 9x3 - 56x2 + 290x - 5236 f(x) = x4 - 9x3 + 56x2 - 290x + 5236 f(x) = x4 - 145x2 + 580x - 5236 f(x) = x4 - 25x2 + 580x - 5236 these are my options @precal

OpenStudy (precal):

ok just doing my calculus homework

OpenStudy (precal):

sorry gotta distracted

OpenStudy (precal):

(x-7)(x+11)(x-(2+8i))(x+(2+8i))

OpenStudy (precal):

you need to multiply out the (x-(2+8i))(x+(2+8i)) first recall i^2 is -1